OpenStudy (anonymous):

helllp plz A falcon flew 70 miles in 3 hours. What was the speed of the falcon? speed = distance over time; 1 mile = 1609 meters.

OpenStudy (anonymous):

speed = 70/3 = 25.3 miles per hour or 70*1609/3=37543 meters per hour

OpenStudy (anonymous):

how do i find the speed of the falcom?